# Retrieve Anyone: A General-purpose Person Re-identification Task with Instructions

### v1

@@ 4.1 Instruction Generation
In our proposed instruct-ReID task, the model must retrieve the images that describe the same person following the instructions. By designing different instructions, our instruct-ReID can be specialized to existing ReID tasks, i.e., traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, and language-instructed ReID. We show the current instructions below and leave the exploration of better instructions for instruct-ReID to future research.
Traditional ReID: The instruction is the sentence "Do not change clothes". The model is expected to retrieve images of the same person without altering image attributes, such as clothing. ### Query image:{Query image}  ### Instruction: "Do not change clothes."
@@
Clothes-changing ReID: The instruction is the sentence "Ignore clothes". The model should retrieve images of the same person even when wearing different outfits. ### Query image:{Query image}  ### Instruction: "Ignore clothes."
@@
Clothes template based clothes-changing ReID: The instruction is a clothes template for a query image while a cropped clothes image for a target image. The model should retrieve images of the same person wearing the provided clothes. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}  ### Instruction:{Any clothes template image}
@@
Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. We randomly select the description languages from the person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person described in the provided sentences. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}  ### Instruction:{Sentences describing pedestrians}
@@ 4.2 Editing Transformer
We illustrate the detailed structure of each layer in the editing transformer in Fig. 3 (right). Given the features $\mathbf{F}_{l-1}$ and instruction features $\mathbf{F}_{\mathbf{T}}$, the attention map $\mathbf{M}_{l}$ is defined as
@@
where $\text{Linear}_{o}$ is the feed-forward network after the attention layer in each transformer block, $\mathbf{V}_{l}$ and $\mathbf{V}^{\prime}_{l}$ are the values calculated by $\mathbf{V}_{l}=\text{Linear}_{v}(\mathbf{F}_{l-1})$ and $\mathbf{V}^{\prime}_{l}=\text{Linear}_{v^{\prime}}(\mathbf{F}_{\mathbf{T}})$, respectively. We use the [CLS] token in the output feature of $L$-th transformer layer for computing losses and retrieval, i.e., $\mathbf{F}=\mathbf{f}_{L}^{\text{CLS}}$, where $\mathbf{F}_{L}=(\mathbf{f}_{L}^{\text{CLS}},\mathbf{f}_{L}^{1},\mathbf{f}_{L}^{2},...,\mathbf{f}_{L}^{N})$ and $N$ is patch number of query images.
@@ 4.3 Adaptive Triplet Loss
Unlike typical triplet loss that defines positive and negative samples solely based on identities, instruct-ReID requires distinguishing images with different instructions for the same identity. Intuitively, an adaptive margin should be set to push or pull samples based on the instruction difference. Let ($\mathbf{F}^{a}_{i}$, $\mathbf{F}^{r_{1}}_{i}$, $\mathbf{F}^{r_{2}}_{i}$) be the $i$-th triplet in the current mini-batch, where $\mathbf{F}^{a}_{i}$ is an anchor sample, $\mathbf{F}^{r_{1}}_{i}$ and $\mathbf{F}^{r_{2}}_{i}$ are reference samples. We propose a novel adaptive triplet loss as
$\small\mathbf{\mathcal{L}}_{atri}=\frac{1}{N_{tri\dagger}}\sum_{i=1}^{N_{tri\dagger}}\left\{\mathbf{Sign}(\beta_{1}-\beta_{2})\left[d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{1}}_{i})+\left(\beta_{1}-\beta_{2}\right)m-d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{2}}_{i})\right]\right\}_{+}$ (5)
@@
The concept of adaptive triplet loss can be described by Fig. 4(a). We discuss the adaptive loss in two scenarios. In the first scenario, as shown in Fig. 4(b), the margin is set to zero if the triplet has the same identity and the instructions of the two reference samples are equally similar to the instruction of the anchor sample. This makes the distances between the anchor point and the two reference points the same. Second, as shown in Fig. 4(c), when a significant difference in the instruction similarities exists, the margin distance between the anchor and two references turns to be more than the maximum value $m$, forcing the model to learn discriminative features.
Figure 4: Illustration of adaptive triplet loss. Unlike traditional triplet loss where the margin is fixed, the margin in our adaptive triplet loss is defined by the instruction similarity for the two query-instruction pairs that describe the same person. The features associated with similar instructions are pulled to be close to each other.
@@ 5.1 Implementation Details
Training Setups. For the editing transformer, we use the plain ViT-Base with the supervised ImageNet pretraining [38]. The instruction encoder is a fixed FashionCLIP [5]. All images are resized into 256$\times$128 for training and testing. We use the SGD optimizer with a base learning rate of 1e-3 and a weight decay of 5e-4 . We linearly warmup the learning rate from 1e-5 to 1e-3 for the first 1500 iteration steps. For STL scenario, only one V100 GPU is used to train the network with 24000 iterations. For MTL scenario, eight V100 GPUs are used. To deal with multi-task distributed training, following the practice in [37; 7], each GPU is responsible for one dataset during training. Random cropping, flipping, and erasing are used for data augmentation during training. For each training batch, we randomly select 32 identities with 4 image samples for each identity.
@@ 5.2 Experimental Results
Clothes-Changing ReID (CC-ReID). As observed in Table 2, RAM outperforms all state-of-the-art methods on LTCC, PRCC and VC-Clothes, showing that the model can effectively extract clothes-invariant features following the "Ignore clothes" instruction. Specifically, on STL, RAM improves the performance of CAL[11], TransReID [12] by +8.9% mAP and +2.9% mAP on LTCC and VC-Clothes, respectively. Using multi-task training further improves the performance of RAM to 56.1% mAP on LTCC, and to reach a new state-of-the-art on PRCC with a 48.0% mAP. While multi-task learning leads to slightly lower performance, our method still achieves a higher Top-1 than TransReID. We conjecture that the performance drop is due to the domain gap between VC-Clothes (Synthetic) and other datasets (Real) and leave it for feature work.
Table 2: The performance of Clothse-Changing ReID of our method and the state-of-the-art methods. Mean average precison (mAP) and Top1 are used to quantify the accuracy.
Methods Modality LTCC PRCC VC-Clothes
@@
HACNN[26] RGB 26.7 60.2 - 21.8 - -
RGA-SC[47] RGB 27.5 65.0 - 42.3 67.4 71.1
PCB[36] RGB 30.6 65.1 38.7 41.8 62.2 62.0
IANet[14] RGB 31.0 63.7 45.9 46.3 - -
CAL[11] RGB 40.8 74.2 - - - -
TransReID[12] RGB - - - 44.2 71.8 72.0
RAM (STL) RGB 49.7 89.0 41.5 41.5 74.7 86.7
RAM(MTL) RGB 56.1 92.9 48.0 47.8 72.0 84.9
Clothes Template Based Clothes-Changing ReID (CTCC-ReID). Our method achieves desirable performance on the CTCC-ReID task in Tabel 3, which shows that a fixed instruction encoder is enough for this tough task. Concretely, when only trained on COCAS+ Real1, RAM outperforms BC-Net [46] and DualBCT-Net [23], both of which learn an independent clothes branch, by +8.6% mAP and +1.2% mAP, respectively. By integrating the knowledge on other instruct ReID tasks during multi-task learning, we are able to further improve the performance of RAM, achieving an mAP of 42.5% and pushing the performance limits on CTCC-ReID.
Table 3: Performance comparison with the state-of-the-art methods on Clothes -Template Clothes-Changing ReID, Language-Instructed ReID, and Traditional ReID. † denotes that the model is trained with multiple datasets.
@@
HumanBench† [37] - - - - 89.5 - 69.1 - 82.6 -
RAM (STL) 31.2 51.7 31.1 56.9 88.0 99.8 65.6 92.8 85.0 99.1
RAM (MTL)† 42.5 62.9 40.4 60.6 90.0 99.8 69.1 93.7 85.9 99.1
Language-Instructed ReID (LI-ReID). As a new setting in ReID, no previous works have been done to retrieve a person using several sentences as the instruction, therefore, we compare RAM with a straightforward baseline. In the baseline method, only a ViT-Base is trained on COCAS+ Real1 images without utilizing the information from language instruction, leading to poor person re-identification ability. As shown in Table 3, RAM improves the baseline by +16.2% mAP, because RAM can integrate instruction information into identity features. With MTL, RAM achieves extra +9.3% performance gain by using more images and general information in diverse ReID tasks.
Traditional ReID (Trad-ReID). RAM also shows its effectiveness on Traditional ReID tasks in Table 3. Specifically, when trained on a single dataset, compared with SAN [18], RAM achieves comparable performance on Market1501 and +8.6% performance gain on CUHK03. With multi-task training, RAM can further outperform the recent multi-task pretraining SOTA HumanBench [37].
@@ 5.3 Ablation Study
Table 4: Ablation study. The performance comparisons between fusion methods, and the comparisons with triplet loss and the proposed adaptive triplet loss in terms of mAP.
Methods CTCC-ReID LI-ReID CC-ReID Trad-ReID Avg.
Real2 Real2 LTCC PRCC VC-Clothes Market1501 MSMT17 CUHK03
RAM (STL) 31.2 31.1 49.7 41.5 74.7 88.0 66.9 85.0 58.5
Late Fusion 31.3 30.2 49.0 41.3 73.1 87.4 65.6 84.5 57.8
Triplet loss 30.1 30.1 49.7 41.5 74.7 88.0 66.9 85.0 58.3
RAM (MTL) 42.5 40.4 56.1 48.0 72.0 90.0 69.1 85.9 63.0
Late Fusion 41.4 37.6 55.6 47.3 71.6 89.6 67.6 85.0 62.0
Triplet loss 42.2 39.7 55.0 47.0 71.0 89.1 66.8 85.0 62.0
@@
Editing Transformer. To verify the effectiveness of the instruction integrating design in the editing transformer, we compare it with Late Fusion, where dual cross-attentions are used between tokens from different models/modalities to fuse information. Results in Table 4 show that adopting Late Fusion leads to -0.7% mAP performance drop in the STL scenario. Consistent results can be observed in MTL, inidicating the effectiveness of the instruction integrating design in the editing transformer.
@@
Adaptive Triplet Loss. Table 4 shows that adaptive triplet loss in MTL outperforms the traditional triplet loss by +1.0% mAP on average, indicating that the proposed loss boosts the model to learn more discriminative features following different instructions. In the STL scenario, for CC-ReID and Trad-ReID, the instructions are fixed sentences leading to the same performance of adaptive/traditional triplet loss. However, in the case of CTCC-ReID and LI-ReID, where instructions vary among samples, using adaptive triplet loss brings about +1.0% mAP performance gain, which demonstrates the effectiveness of adaptive triplet loss in learning both identity and instruction similarity.
6 Visualizaton
We visualize the retrieval results of CTCC-ReID, LI-ReID, CC-ReID and Trad-ReID tasks in Fig. 5. Given a query image, RAM not only retrieves the right person from the gallery but also finds specific target images of the person following the instruction. Concretely, for CTCC-ReID, RAM retrieves images of query persons wearing instructed clothes as shown in the first row. For LI-ReID, RAM effectively parses information from languages such as bag condition (e.g., row 2 person #1) and clothes attribute (e.g., row 2 person #2,3) to find the correct image. For CC-ReID, with the "Ignore clothes" instruction, our method focuses on biometric features and successfully retrieves the person in the case of changing clothes, e.g. the 4th and 5th images of row 3 person #1, which the clothes are different from the query image. For Trad-ReID shown in row 4, "do not change clothes" instructs RAM to pay attention to clothes, a main feature of a person’s image. In this case, RAM retrieves images with the same clothes in the query image.
Figure 5: Illustration of all tasks retrieval results by RAM. We visualize the task-specific instructions on three people for example. Green and red boxes mean true and false matches.
7 Conclusion and Limitations

### v4

@@ 4.1 Instruction Generation
In our proposed instruct-ReID task, the model must retrieve the images that describe the same person following the instructions. By designing different instructions, our instruct-ReID can be specialized to existing ReID tasks, i.e., traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, visual-infrared ReID, text-to-image ReID, and language-instructed ReID. We show the current instructions and leave the exploration of better instructions for instruct-ReID to future research.
Traditional ReID: Following Instruct-BLIP [8], we generate 20 instructions [^3] See all available instruction sentences in supplementary materials. from GPT-4 and randomly select one, e.g., “Do not change clothes.”, when training. The model is expected to retrieve images of the same person without altering image attributes, such as clothing. ### Query image:{Query image}
### Instruction: “Do not change clothes.”${}^{3}$
@@
[…]
### Instruction: “With clothes changed.”${}^{3}$
@@
Clothes template based clothes-changing ReID: The instruction is a clothes template for a query image while a cropped clothes image for a target image. The model should retrieve images of the same person wearing the provided clothes. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}
### Instruction:{Any clothes template image}
@@
[…]
### Instruction: “Retrieve cross modality image.”${}^{3}$
### Target image:{Output}
Text-to-Image ReID: The instruction is the describing sentences, and both images and text are fed into IRM during the training process. While in the inference stage, the image features and instruction features are extracted separately.  [^4] Image and instruction features are extracted separately in test stage. ### Image:{Image}${}^{4}$
### Instruction:{Sentences describing pedestrians}${}^{4}$
### Target image:{Output}
Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. We randomly select the description languages from the person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person described in the provided sentences. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}
### Instruction:{Sentences describing pedestrians}
@@ 4.2 Editing Transformer
We show the detailed structure of each layer in the editing transformer in Fig. 3 (b). Given the features $\mathbf{F}_{l-1}$ and instruction features $\mathbf{F}_{\mathbf{T}}$, the attention map $\mathbf{M}_{l}$ is defined as
@@
where $\text{Linear}_{o}$ is the feed-forward network after the attention layer in each transformer block, $\mathbf{V}_{l}$ and $\mathbf{V}^{\prime}_{l}$ are the values calculated by $\mathbf{V}_{l}=\text{Linear}_{v}(\mathbf{F}_{l-1})$ and $\mathbf{V}^{\prime}_{l}=\text{Linear}_{v^{\prime}}(\mathbf{F}_{\mathbf{T}})$. We use the [CLS] token in the output feature of $L$-th transformer layer for computing losses and retrieval, i.e., $\mathbf{F}=\mathbf{f}_{L}^{\text{CLS}}$, where $\mathbf{F}_{L}=(\mathbf{f}_{L}^{\text{CLS}},\mathbf{f}_{L}^{1},\mathbf{f}_{L}^{2},...,\mathbf{f}_{L}^{N})$ and $N$ is patch number of query images.
@@ 4.3 Adaptive Triplet Loss
Unlike typical triplet loss that defines positive and negative samples solely based on identities, instruct-ReID requires distinguishing images with different instructions for the same identity. Intuitively, an adaptive margin should be set to push or pull samples based on the instruction difference. Let ($\mathbf{F}^{a}_{i}$, $\mathbf{F}^{r_{1}}_{i}$, $\mathbf{F}^{r_{2}}_{i}$) be the $i$-th triplet in the current mini-batch, where $\mathbf{F}^{a}_{i}$ is an anchor sample, $\mathbf{F}^{r_{1}}_{i}$ and $\mathbf{F}^{r_{2}}_{i}$ are reference samples. We propose an adaptive triplet loss as
$\small\begin{split}\mathbf{\mathcal{L}}_{atri}=\frac{1}{N_{tri\dagger}}\sum_{i=1}^{N_{tri\dagger}}\{\mathbf{Sign}(\beta_{1}-\beta_{2})[d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{1}}_{i})+\\ \left(\beta_{1}-\beta_{2}\right)m-d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{2}}_{i})]\}_{+}\end{split}$ (5)
@@
The concept of adaptive triplet loss is described by Fig. 4(a). We discuss the adaptive loss in two cases. First, as shown in Fig. 4(b), the margin is set to zero if the triplet has the same identity and the instructions of the two reference samples are equally similar to the instruction of the anchor sample. This makes the distances between the anchor point and the two reference points the same. Second, as shown in Fig. 4(c), when there is a significant difference in instruction similarities, the margin distance between the anchor and two references becomes closer to the maximum value $m$, forcing the model to learn discriminative features. Adaptive triplet loss makes features from the same person become distinctive based on the similarity of instructions, which helps to retrieve images that align the requirements of given instructions in the CTCC-ReID and LI-ReID tasks.
Figure 4: Illustration of adaptive triplet loss. Unlike the traditional triplet loss where the margin is fixed, the margin in our adaptive triplet loss is defined by the instruction similarity for the two query-instruction pairs that describe the same person. The features associated with similar instructions are pulled to be closer.
4.4 Overall Loss Function
We impose an identification loss $\mathcal{L}_{id}$, which is the classification loss on identities, and an adaptive triplet loss $\mathcal{L}_{atri}$ on $\mathbf{F}$ and the final fusion features $\mathbf{F}_{\mathbf{out}}$ to supervise the model for training Trad-ReID, CC-ReID, VI-ReID, CTCC-ReID, and LI-ReID tasks. The overall loss is combined as
$\small\begin{split}\mathbf{\mathcal{L}}=\mathcal{L}_{atri}(F)+\mathcal{L}_{id}(F)+\mathcal{L}_{atri}(F_{out})+\mathcal{L}_{id}(F_{out})\end{split}$ (7)
where $\mathbf{F}$ and $\mathbf{F}_{out}$ indicate the source of features used in calculating the loss.
For the T2I-ReID task, we adopt a contrastive loss $\mathcal{L}_{cl}$ to align the image features $\mathbf{F}$ and text features $\mathbf{F}_{\mathbf{T}}$ to enable text-based person retrieval. We also employ a binary classification loss $\mathcal{L}_{match}$ to learn whether an inputted image-text pair is positive or negative, defined as:
$\small\begin{split}\mathbf{\mathcal{L}}=\mathcal{L}_{cl}({F})+\mathcal{L}_{match}({F_{out}})\end{split}$ (8)
where, $\mathcal{L}_{match}({F_{out}})=\mathcal{C}_{e}(\hat{y},F_{out})=\mathcal{C}_{e}[\hat{y},\mathcal{E}_{f}(\mathbf{F},\mathbf{F_{T}})]$, respectively. $\mathcal{C}_{e}$ represents a binary cross-entropy function, $\hat{y}$ is a 2-dimension one-hot vector representing the ground-truth label i.e., $[0,1]^{\intercal}$ for the positive pair and $[1,0]^{\intercal}$ for the negative pair, which is formed by matching text features with corresponding image features before inputting into the attention module $\mathcal{E}_{f}$.
@@ 5.1 Experimental Setups
Implementation Details. For the editing transformer, we use the plain ViT-Base with the ReID-specific pretraining [69]. The instruction encoder is ALBEF [22]. All images are resized into 256$\times$128 for training and test. We use the AdamW optimizer with a base learning rate of 1e-5 and a weight decay of 5e-4 . We linearly warmup the learning rate from 1e-7 to 1e-5 for the first 1000 iterations. Random cropping, flipping, and erasing are used for data augmentation during training. For each training batch, we randomly select 32 identities with 4 image samples for each identity.
Table 2: The performance of Clothse-Changing ReID of our method and the state-of-the-art methods. Mean average precision (mAP) and Top1 are used to quantify the accuracy. † denotes that the model is trained with multiple datasets. * denotes that the model is pre-trained on 4 million images.
Methods LTCC PRCC VC-Clothes
mAP Top1 mAP Top1 mAP Top1
HACNN[29] 26.7 60.2 - 21.8 - -
RGA-SC[63] 27.5 65.0 - 42.3 67.4 71.1
PCB[43] 30.6 65.1 38.7 41.8 62.2 62.0
IANet[16] 31.0 63.7 45.9 46.3 - -
CAL[12] 40.8 74.2 - - - -
TransReID[14] - - - 44.2 71.8 72.0
IRM (STL) 46.7 66.7 46.0 48.1 80.1 90.1
IRM (MTL)† 52.0 75.8 52.3 54.2 78.9 89.7
@@ 5.2 Experimental Results
Clothes-Changing ReID (CC-ReID). As shown in Tab 2, IRM outperforms all state-of-the-art methods on LTCC, PRCC and VC-Clothes, showing that the model can effectively extract clothes-invariant features following the instructions, e.g., “Ignore clothes”. Specifically, on STL, IRM improves CAL [12], TransReID [14] by +5.9% mAP and +8.3% mAP on LTCC and VC-Clothes, respectively. Using multi-task training further improves the performance of IRM to 52.0% mAP on LTCC, and to reach a new state-of-the-art on PRCC with a 52.3% mAP. While multi-task learning leads to slightly lower performance than single-task learning on VC-Clothes, our method still achieves a higher Top-1 than TransReID. We conjecture that this drop is due to the domain gap between VC-Clothes (Synthetic) and other datasets (Real) and leave it for future work.
Clothes Template Based Clothes-Changing ReID (CTCC-ReID). Our method achieves desirable performance on the CTCC-ReID task in Tabel 4, which shows that a fixed instruction encoder is enough for this tough task. Concretely, when only trained on COCAS+ Real1, IRM outperforms BC-Net [57] and DualBCT-Net [28], both of which learn an independent clothes branch, by +9.6% mAP and +2.2% mAP, respectively. By integrating the knowledge on other instruct ReID tasks during multi-task learning, we are able to further improve the performance of IRM, achieving an mAP of 41.7% and pushing the performance limits on CTCC-ReID.
Table 3: Comparison with the state-of-the-art methods on visible-infrared ReID and text-to-image ReID. The VI-ReID setting is VIS-to-IR and IR-to-VIS in LLCM. † denotes that the model is trained with multiple datasets.* denotes that the model is trained under the same image shape as IRM, i.e., 256$\times$128.
Methods T2I-ReID VI-ReID: LLCM
CUHK-PEDES VIS-to-IR IR-to-VIS
@@
ALBEF [22] 56.7 60.3 - - - -
CAJ [54] - - 59.8 56.5 56.6 48.8
MMN [62] - - 62.7 59.9 58.9 52.5
DEEN [61] - - 65.8 62.5 62.9 54.9
SAF [27] 58.6 64.1 - - - -
PSLD [13] 60.1 64.1 - - - -
RaSa* [3] 63.9 69.6 - - - -
IRM (STL) 65.3 72.8 66.6 66.2 64.5 64.9
IRM (MTL)† 66.5 74.2 67.5 66.7 67.2 65.7
Table 4: Performance comparison with the state-of-the-art methods on Clothes -Template Clothes-Changing ReID, Language-Instructed ReID, and Traditional ReID. † denotes that the model is trained with multiple datasets. * denotes that the model is pretrained on 4 million pedestrian images. The size of the input images used in the table is 256x128.
@@
HumanBench† [44] - - - - 89.5 - 69.1 - 77.7 -
PASS* [69] - - - - 93.0 96.8 71.8 88.2 - -
IRM (STL) 32.2 54.8 30.7 60.8 92.3 96.2 71.9 86.2 83.3 86.5
IRM (MTL)† 41.7 64.9 39.8 71.6 93.5 96.5 72.4 86.9 85.4 86.5
Visible-Infrared ReID (VI-ReID). We validate the performance of IRM on Visible-Infrared ReID datasets LLCM, which is a new and challenging low-light cross-modality dataset and has a more significant number of identities and images. From Tab. 3, we can see that the results on the two test modes show that the proposed IRM achieves competitive performance against all other state-of-the-art methods. Specifically, for the VIS-to-IR mode on LLCM, IRM achieves 67.5% mAP and exceeds previous state-of-the-art methods like DEEN [61] by +1.7%. For the IR-to-VIS mode on LLCM, IRM achieves 65.7% Rank-1 accuracy and 67.2% mAP, which is a new state-of-the-art result. The results validate the effectiveness of our method.
Text-to-Image ReID (T2I-ReID). As shown in Tab. 3, IRM shows competitive performance with a mAP of 66.5% on the CUHK-PEDES [26], which is +2.6%, +6.4%, +7.9% higher than previous methods RaSa [3] (63.9%), PSLD [13] (60.1%), SAF [27] (58.6%). Because a few images in CUHK-PEDES training set are from the test sets of Market1501 and CUHK03, we filtered out duplicate images from the test sets during the multi-task learning (MTL) process. The testing was conducted on uniformly resized images with a resolution of 256$\times$128.
Table 5: Ablation study. The performance comparison of our editing transformer and using ViT base transformer (w/o editing), and the comparisons with triplet loss (w/o $\mathcal{L}_{atri}$) and the proposed adaptive triplet loss in terms of mAP. † denotes that the test mode is VIS-to-IR and ‡ denotes IR-to-VIS mode on LLCM.
Methods CTCC-ReID LI-ReID T2I-ReID VI-ReID CC-ReID Trad-ReID Avg.
Real2 Real2 CUHK. LLCM† LLCM‡ LTCC PRCC VC-Clo. Market1501 MSMT17 CUHK03
IRM (STL) 32.2 30.7 65.3 66.6 64.5 46.7 46.0 80.1 92.3 71.9 83.3 61.8
w/o editing 32.6 30.5 65.7 66.2 65.1 46.2 45.7 77.8 92.5 71.4 83.2 61.5
w/o $\mathcal{L}_{atri}$ 31.5 30.2 - 66.6 64.5 46.7 46.0 80.1 92.3 71.9 83.3 61.3
IRM (MTL) 41.7 39.8 66.5 67.5 67.2 52.0 52.3 78.9 93.5 72.4 85.4 65.1
w/o editing 41.0 38.8 66.1 67.3 65.2 51.2 51.0 78.6 93.0 71.8 85.1 64.5
w/o $\mathcal{L}_{atri}$ 40.7 39.2 65.2 68.4 66.3 52.0 52.4 77.9 92.9 72.0 85.5 64.8
Language-Instructed ReID (LI-ReID). As a new setting in ReID, no previous works have been done to retrieve a person using several sentences as the instruction, therefore, we compare IRM with a straightforward baseline. In the baseline method, only a ViT-Base is trained on COCAS+ Real1 images without utilizing the information from language instruction, leading to poor person re-identification ability. As shown in Tab 4, IRM improves the baseline by +15.8% mAP, because IRM can integrate instruction information into identity features. With MTL, IRM achieves extra +9.1% performance gain by using more images and general information in diverse ReID tasks.
Traditional ReID (Trad-ReID). IRM also shows its effectiveness on Trad-ReID tasks in Tab 4. Specifically, when trained on a single dataset, compared with PASS [69], IRM achieves comparable performance on Market1501, MSMT17 and +6.9% performance gain on CUHK03 compared with SAN [19]. With multi-task training, IRM can outperform the recent multi-task pretraining HumanBench [44] and self-supervised pretraining [69]. We do not compare with SOLDIER [5] because it only reports ReID results with the image size of 384x192 instead of 256x128 in our method. More importantly, SOLDIER primarily focuses on pretraining and is evaluated on traditional ReID tasks only, while the claimed contribution of IRM is to tackle multiple ReID tasks with one model.
@@ 5.3 Ablation Study
Editing Transformer. To verify the effectiveness of the instruction integrating design in the editing transformer, we compare it with the traditional ViT base model, where the image features are extracted without fusing information from instruction features. Results in Tab 5 show that adopting the editing transformer leads to -0.6% mAP performance drop in the MTL scenario. Consistent results can be observed in STL, indicating the effectiveness of the instruction integrating design in the editing transformer.
@@
Adaptive Triplet Loss. Tab 5 shows that adaptive triplet loss in MTL outperforms the traditional triplet loss by +0.3% mAP on average, indicating that the proposed loss boosts the model to learn more discriminative features following different instructions. In the STL scenario, for CC-ReID and Trad-ReID, the instructions are fixed sentences leading to the same performance of adaptive/traditional triplet loss. However, in the case of CTCC-ReID and LI-ReID, where instructions vary among samples, using adaptive triplet loss brings about +0.7%, +0.5% mAP gain, which shows the effectiveness of adaptive triplet loss in learning both identity and instruction similarity.
More Results. We provide additional results based on different pre-trained models and visualizations of the retrieval results in the supplementary material.
6 Conclusion